# AutoSAM: Towards Automatic Sampling of User Behaviors for Sequential Recommender Systems

### v3

@@ Abstract.
Sequential recommender systems (SRS) have garnered immense popularity in recommendation due to their remarkable proficiency in capturing dynamic user preferences. In the current setup of SRS, a common configuration is to uniformly consider each historical behavior as a positive interaction. However, this setting has the potential to yield sub-optimal performance as each individual item often have a different impact on shaping the user’s interests. For instance, purchased items should hold more significance than merely clicked ones. Hence, in this paper, we propose a novel automatic sampling framework for sequential recommendation, named AutoSAM, to non-uniformly treat historical behaviors. Specifically, AutoSAM extends the conventional SRS framework by integrating an extra sampler to intelligently discern the skew distribution of the raw input, and then sample informative sub-sets to build more generalizable SRS. To tackle the challenges posed by non-differentiable sampling actions and to introduce multiple decision factors for sampling, we further design a novel reinforcement learning based method to guide the training of the sampler. We theoretically devise multi-objective sampling rewards including Future Prediction and Sequence Perplexity, and then optimize the whole framework in an end-to-end manner by combining the policy gradient. We conduct a broad range of tests on benchmark recommendation models and four real-world datasets from different online platforms. The experimental results demonstrate the effectiveness of the proposed approach. We also make our code publicly available  [^1] https://github.com/zh-ustc/AutoSAM.
Sequential Recommendation, User Interest, Automatic Sampling
Qi Liu is corresponding author.
conference: XXX; XXX; ccs: Information systems Recommender System
1. Introduction
@@
Recommender systems  (Koren, 2008; Koren et al., 2009; Liu et al., 2018; Wu et al., 2022) have become crucial tools for information filtering in various online applications, such as e-commerce, advertising, and online videos. Among them, sequential recommender systems (SRS)  (Huang et al., 2018; Wang et al., 2019; Li et al., 2017; Liu et al., 2022; Singer et al., 2022) have become increasingly prevalent due to their ability to capture long-term and short-term user interests.
@@
So far, many efforts have been devoted to sequential recommendation, ranging from early matrix factorization and Markov chain based method (Rendle et al., 2010) to state-of-the-art deep neural network models, including recurrent neural networks (RNNs) (Hidasi et al., 2015, 2016), convolutional neural networks (CNNs) (Tang and Wang, 2018; Yuan et al., 2019), and self-attentive models (Kang and McAuley, 2018; Sun et al., 2019). Meanwhile, the effectiveness and efficiency of the generative loss (i.e., auto-regressive loss) (Yuan et al., 2019) in SRS tasks have been widely demonstrated in these approaches. Despite their remarkable success, these methods often consider from a model perspective while treating all historical behaviors as uniformly positive. In fact, this may lead to sub-optimal performance as items in the sequences are usually of unequal importance on shaping user interests (Zhang et al., 2013). That is because that the motivation behind the user behaviors are often very complex and multifaceted. For example, some interactions on online platforms may be driven by a herd mentality, specific purchasing needs or even the result of mis-clicks, while others may represent the user’s true preferences.
@@
Recently, there have also been some works which attempt to improve the recommender systems from a data perspective (He et al., 2023; Qin et al., 2023). To be specific, SIM (Pi et al., 2020) extracts user interests with two cascaded search units to capture the diverse user’s long-term interest with target item. Similarly, UBR4CTR (Qin et al., 2020) adopts two-stage frameworks to model long-term interests. At the first stage, it generates the query according to the target item for retrieving similar historical behaviors. Then, at the second stage, these items are leveraged to predict the user’s interests. To reduce the information loss of the above methods, SDIM (Cao et al., 2022) proposes a sampling-based end-to-end approach for CTR prediction, which samples from multiple hash functions to gather similar behavior items to the target. Though the aforementioned methods have proven effective, they may suffer from the following limitations. Firstly, these methods often select items according to the target, which brings challenges to compute scores across all candidate items in parallel and may be only suitable to fine-ranking stage (Covington et al., 2016). Besides, such a setup may be too strict and fail to consider from a dynamic perspective. Secondly, these sampling processes are mainly designed for CTR prediction tasks (Guo et al., 2017; Wang et al., 2017), which are often typically optimized as binary classification problems. As a result, these approaches often cannot be equipped with generative loss in SRS tasks, whose effectiveness and efficiency have been demonstrated in many previous works (Kang and McAuley, 2018; Yuan et al., 2019).
We hold that an ideal solution should be adaptive while maintaining the core training principles of SRS. Hence, to achieve this goal, we propose a general automatic sampling framework, named AutoSAM, to non-uniformly treat historical behaviors for sequential recommendation. To be concrete, an additional sampler layer is first employed to adaptively explore the skew distribution of raw input. Then, as shown in Figure 1, informative sub-sets are sampled from the distribution to build more generalizable sequential recommenders. Meanwhile, the removal of some low-signal items releases the model from the constraint of always taking the next step as the target in traditional SRS, which helps explore more potential preferred items in the future.
Figure 1. An illustration of a user’s generated samples, items outlined by red boxes stand for the user’s favorite, while the green box means misclicks, etc.
@@
In order to overcome the challenges of non-differentiable discrete sampling actions, and to introduce multiple decision factors for sampling, we further introduce a novel reinforcement learning based method to guide the training of the sampler. We believe that a proper decision should not only focus on the target item in the future, but also consider the coherence with the previous context. Along this line, we incorporate the major factors into the reward estimation, including Future Prediction and Sequence Perplexity. Finally, both the sequential recommender and the sampler can be jointly optimized in an end-to-end manner by combining the policy gradient. As a result, our method exhibits more accurate recommendations than previous approaches, and is generally effective for various sequential recommenders with different backbones. We summarize the contributions as follows:
@@
    We propose to sample user behaviors from a non-uniform distribution for sequential recommendation. We highlight that the main challenge is to sample historical behaviors dynamically while leveraging the benefits of generative loss.
@@
    We propose a general automatic sampling framework for sequential recommendation, named as AutoSAM, to build generalizable SRS with an additional sample layer. We further introduce a reinforcement learning based method to solve the challenges of non-differentiable decisions and design multi-objective rewards to guide the training of the sampler.
@@
Figure 2. (a) is the architecture of the standard sequential recommender system (SRS). (b) is our proposed AutoSAM which augments the traditional sequential recommendation architecture with an additional sampler layer to adaptively explore the skew distribution of raw input. The dashed lines represent gradient backpropagation.
2. Preliminaries
Before going into the details, we first introduce the research problem and revisit the sequential recommender systems.
2.1. Problem Definition
@@
2.2. Sequential Recommender System
As shown in Figure 2(a), modern SRS mainly contain the components as following. First, each input item $x^{u}_{t}$ is mapped into an embedding vector by an embedding lookup operation as:
(1) $E_{t}=\zeta_{x_{t}},$
where $\zeta\in\mathbb{R}^{|\mathcal{I}|\times d}$ is item embedding matrix. Then, the sequential embedding will be fed into stacked hidden layers in order to capture the long- and short-term dependence.
(2) $H^{(l)}$ $=F^{(l)}(H^{(l-1)}),1\leq l\leq L,$
where $L$ stands for the number of layers, $H^{(l)}$ is the output of the $l$-th hidden layer $F^{(l)}$ while $H^{(0)}=E$. Next, a prediction layer is adopted to generate the distribution $P$ of the user’s next preferences for each time step $t$, which can be written as:
(3) $P_{t}=Softmax(H^{(L)}_{t}W+b),$
in which $W\in\mathbb{R}^{d\times|\mathcal{I}|}$, $b\in\mathbb{R}^{|\mathcal{I}|}$ is the learnable parameters. Finally, to train the model, the auto-regressive generative loss are employed as optimization objectives, i.e., left-to-right supervision signals. Combining the cross-entropy loss with such an objective, the loss function can be written as follows:
(4) $\mathcal{L}(X;\theta)=-\sum_{t=1}^{n-1}logP(y_{t}|X_{\leq t};\theta)=-\sum_{t=1}^{n-1}logP_{t,y_{t}},$
where $\theta$ denotes the parameters of the model, $y_{t}=x_{t+1}$ is the item expect to be predicted at step $t$, and $n$ is the input length.
3. AutoSAM: the Proposed Method
3.1. Overview of the Framework
@@
3.2. User Behavior Sampler
@@
(6) $\pi=Softmax(S/\tau)\in\mathbb{R}^{n\times 2}\;\;\;\;\;\;\;\;\;\;$
$A\sim\pi\in\{0,1\}^{n},$
@@
3.3. Multi-Objective Rewards
Now we discuss how to design our reward, which plays an important role in learning the optimal $\pi$. As mentioned before, traditional SRS often follow the next item prediction task, which can neither update sampler network directly due to the non-differentiable challenge of discrete actions nor control the sample rate. In addition, such single objective may be too strict and one-sided. We argue that the proper decisions should not only focus on the target item but also consider the coherence with preceding context. Along this line, in this paper, we incorporate the major factors into the reward estimation, including Future Prediction and Sequence Perplexity.
3.3.1. Future Prediction
@@
(7) $min\;J(\theta,\phi)$ $=min\;\mathbb{E}_{X}\mathbb{E}_{A\sim\pi_{\phi}(A|X)}\left[\mathcal{L}(\hat{X};\theta)\right],$
where $\theta$ and $\phi$ are the parameters of SRS and sampler, respectively, while sharing the embedding layer. To optimize this objective, we first transform the prediction loss to an equivalent form:
(8) $\mathcal{L}(\hat{X};\theta)$ $=\sum_{t=1}^{m-1}-logP(\hat{y}_{t}|\hat{X}_{\leq t};\theta)$
@@
(9) $\nabla_{\phi}J(\theta,\phi)$ $=\mathbb{E}_{X}\nabla_{\phi}\sum_{A}\pi_{\phi}(A|X)\mathcal{L}(\hat{X};\theta)$
@@
where $r^{pre}=\sum_{t=1}^{n-1}-\mathcal{L}_{t}(X,A;\theta)$ means the rewards associated to the sampler. More details of mathematical derivations can be found in Appendix A.1. In practice, we may make a few changes to it. First, we add a baseline $\mathcal{L}^{b}$ computed on raw input to $\mathcal{L}$ as it is always positive. Second, we discount the future rewards by a factor $\gamma\in[0,1]$ and ignore the targets whose sample probabilities are lower than $\psi$ to increase stability. Then, the future prediction reward at each time step $t$ can be defined as follows:
(10) $r^{pre}_{t}=\sum_{t^{\prime}=t}^{n-1}-\gamma^{t^{\prime}-t}I[\pi_{\phi}(a_{t^{\prime}+1}|X)>\psi](\mathcal{L}_{t^{\prime}}(X,A;\theta)-\mathcal{L}_{t^{\prime}}^{b}),$
@@
3.3.2. Sequence Perplexity
In addition to sampling based on future predictions, it should also be attached great importance to considering the coherence with previous context. Inspired by perplexity (PP), which is widely used to evaluate the quality of sentences in Natural Language Processing (NLP) (Liu et al., 2016; Dathathri et al., 2019) with the definition as follows:
(11) $PP(X)=P(x_{1},x_{2},...x_{n})^{-\frac{1}{n}}=\prod_{t=1}^{n}P(x_{t}|X_{<t})^{-\frac{1}{n}}.$
As words causing high perplexity tend to stray from the context, similarly, we insist that the items causing high perplexity may lack representativeness. A natural idea is to assign lower sampling probabilities to the items whose predicted probabilities according to historical behaviors are lower than $1/PP(X)$, so as to reduce the perplexity and obtain informative sub-sets in some way. To this end, the sampler should be encouraged to drop $x_{t}$ ($a_{t}=0$) while $P(x_{t}|X_{<t})<1/PP(X)$, and select $x_{t}$ while $P(x_{t}|X_{<t})>1/PP(X)$.
Note that the perplexity can be transformed to cross-entropy, in practice, we employ the SRS to compute the approximate objective at each time step $t$. Besides, a relax factor is further conducted to control the strictness of the sampler. Above all, the reward can be defined as follow:
(12) $r^{pp}_{t}=\left(logP(x_{t}|X_{<t})-log\frac{1}{PP(X)}+b\right)(2a_{t}-1),$

### v4

# Towards Automatic Sampling of User Behaviors for Sequential Recommender Systems

@@ Abstract
Sequential recommender systems (SRS) have gained increasing popularity due to their remarkable proficiency in capturing dynamic user preferences. In the current setup of SRS, a common configuration is to uniformly consider each historical behavior as a positive interaction. However, this setting has the potential to yield sub-optimal performance as each individual item often have a different impact on shaping the user’s interests. Hence, in this paper, we propose a novel automatic sampling framework for sequential recommendation, named AutoSAM, to non-uniformly treat historical behaviors. Specifically, AutoSAM extends the conventional SRS framework by integrating an extra sampler to intelligently discern the skew distribution of the raw input, and then sample informative sub-sets to build more generalizable SRS. To tackle the challenges posed by non-differentiable sampling actions and to introduce multiple decision factors for sampling, we further design a novel reinforcement learning based method to guide the training of the sampler. Furthermore, we theoretically devise multi-objective sampling rewards including Future Prediction and Sequence Perplexity, and then optimize the whole framework in an end-to-end manner by combining the policy gradient. We conduct extensive experiments on benchmark recommendation models and four real-world datasets. The experimental results demonstrate the effectiveness of the proposed AutoSAM [^1] https://github.com/zh-ustc/AutoSAM.
1 Introduction
@@
Recommender systems  (Koren et al., 2009; Liu et al., 2018; Zhang et al., 2025, 2024a) have become crucial tools for information filtering in various online applications, such as e-commerce, advertising, and online videos. Among them, sequential recommender systems (SRS)  (Li et al., 2017; Liu et al., 2022; Singer et al., 2022; Zhang et al., 2024b) have become increasingly prevalent due to their ability to capture long- and short-term user interests Cheng et al. (2021).
@@
So far, many efforts have been devoted to sequential recommendation, ranging from early matrix factorization and Markov chain based method (Rendle et al., 2010) to state-of-the-art deep neural network models Cheng et al. (2022), including recurrent neural networks (RNNs) (Hidasi et al., 2015, 2016), convolutional neural networks (CNNs) (Tang and Wang, 2018; Yuan et al., 2019), and self-attentive models (Kang and McAuley, 2018; Sun et al., 2019; Li et al., 2021). Meanwhile, the effectiveness and efficiency of the generative loss (i.e., auto-regressive loss) (Yuan et al., 2019) in SRS tasks have been widely demonstrated in these approaches. Despite their remarkable success, these methods often consider from a model perspective while treating all historical behaviors as uniformly positive. In fact, this may lead to sub-optimal performance as items in the sequences are usually of unequal importance on shaping user interests (Zhang et al., 2013). For instance, purchased items should hold more significance than merely clicked ones. Moreover, the motivations behind the same action (e.g. clicking) can also be very complex and multifaceted. Interactions on online platforms may sometimes be driven by a herd mentality or could even be the result of mis-clicks, while in other cases, they may accurately represent a user’s genuine interests.
@@
Recently, there have also been some works which attempt to improve the recommender systems from a data perspective (He et al., 2023; Qin et al., 2023). SIM (Pi et al., 2020) and UBR4CTR (Qin et al., 2020) extracts user interests with search or retrieve units to capture the diverse user’s long-term interest with target item. Similarly, SDIM (Cao et al., 2022) samples from multiple hash functions to gather similar behavior items to the target for CTR prediction. Though the aforementioned methods have proven effective, they may suffer from the following limitations. Firstly, these methods often select items according to the target, which brings challenges to compute scores across all candidate items in parallel and may be only suitable to fine-ranking stage (Covington et al., 2016). Besides, such a setup may be too strict and fail to consider from a dynamic perspective. Secondly, these sampling processes are mainly designed for CTR prediction tasks (Guo et al., 2017; Wang et al., 2017), which are often typically optimized as binary classification problems. As a result, these approaches often cannot be equipped with generative loss in SRS tasks, whose effectiveness and efficiency have been demonstrated in many previous works (Kang and McAuley, 2018; Yuan et al., 2019).
@@
We hold that an ideal solution should be adaptive while maintaining the core training principles of SRS. Hence, to achieve this goal, we propose a general automatic sampling framework, named AutoSAM, to non-uniformly treat historical behaviors for sequential recommendation. To be concrete, an additional sampler layer is first employed to adaptively explore the skew distribution of raw input. Then, informative sub-sets are sampled from the distribution to build more generalizable sequential recommenders. In order to overcome the challenges of non-differentiable discrete sampling actions, and to introduce multiple decision factors for sampling, we further introduce a novel reinforcement learning based method to guide the training of the sampler due to its flexibility. We believe that a proper decision should not only focus on the target item in the future, but also consider the coherence with the previous context. Along this line, we incorporate the major factors into the reward estimation, including Future Prediction and Sequence Perplexity. Finally, both the SRS and the sampler can be jointly optimized in an end-to-end manner by combining the policy gradient. As a result, our method exhibits more accurate recommendations than previous approaches, and is generally effective for various sequential recommenders with different backbones. We summarize the contributions as follows:
@@
    We propose to sample user behaviors from a non-uniform distribution for SRS task. We highlight that the main challenge is to sample historical behaviors dynamically while leveraging the benefits of generative loss.
@@
    We propose a general automatic sampling framework for sequential recommendation, named AutoSAM, to build generalizable SRS with an additional sample layer. We further introduce a reinforcement learning based method to solve the challenges of non-differentiable actions and design multi-objective rewards to optimize the sampler.
@@
Figure 1: (a) is the architecture of the standard sequential recommender system (SRS). (b) is our proposed AutoSAM which augments the traditional sequential recommendation architecture with an additional sampler layer to adaptively explore the skew distribution of raw input. The dashed lines represent gradient backpropagation.
2 Related Work
2.0.1 Sequential Recommendation
Sequential recommendation aims to predict users’ future behaviors given their historical interaction data Luo et al. (2024); Cheng et al. (2024). Early approaches mainly fuse Markov Chain and matrix factorization to capture both long- and short-term item-item transitions (Rendle et al., 2010). Latter, with the success of neural network, recurrent neural network (RNN) methods are widely conducted in sequential recommendation (Hidasi et al., 2015, 2016). Besides, convolutional-based models (Tang and Wang, 2018; Yuan et al., 2019) can also be very effective in modeling sequential behaviors. In addition, we notice that graph neural networks (Wu et al., 2019; Xu et al., 2019) have become increasingly prevalent by constructing graph structures from session sequences. And in recent years, self-attention models (Kang and McAuley, 2018; Sun et al., 2019; de Souza Pereira Moreira et al., 2021) have shown their promising strengths in the capacity of long-term dependence modeling and parallel computation. However, these methods focus on the model architectures and treat all behaviors as uniformly positive, may lead to the sub-optimal performance.
2.0.2 User Behavior Sampling
As a data-driven technology, recommendation systems have attracted a series of works that consider from the data perspective in recent years (Qin et al., 2021). To be specific, in SIM (Pi et al., 2020), a two-stage method with a general search unit (GSU) and an exact search unit is proposed to model long-term user behaviors better. Similarly, UBR4CTR (Qin et al., 2020) conducts retrieval-based method to achieve the goal. Besides, SDIM (Cao et al., 2022) proposed a sampling-based approach which samples similar items to the target by multiple hash functions. And some other works (Wang et al., 2021; Lin et al., 2023) mainly formulate a denoising task with the aims of filtering irrelevant or noise items. Despite effectiveness, these methods are mainly designed for CTR prediction, which may be not suitable to conduct the left-to-right generative loss and sample items dynamically in the SRS task. We also notice some works which attempt to capture more informative patterns between user behaviors sequence by designing advance model architectures. RETR (Yao et al., 2022) build recommender transformer for sampling the user behavior pathway. And an all-MLP based model is proposed (Zhou et al., 2022) which adopts Fourier transform with learnable filters to alleviate the influence of the noise item. Differently, in this paper, we present a general framework to learn the distributions of the raw inputs adaptively with carefully designed multi-objectives, so as to enhance the SRS with more informative training data.
3 Preliminaries
3.1 Problem Definition
@@
3.2 Sequential Recommender System
As shown in Figure 1(a), modern SRS contain the components as following. First, each input item $x^{u}_{t}$ is mapped into an embedding vector by an embedding lookup operation as: $E_{t}=\zeta_{x_{t}},$ where $\zeta\in\mathbb{R}^{|\mathcal{I}|\times d}$ is item embedding matrix. Then, the sequential embedding will be fed into stacked hidden layers to capture the long- and short-term dependence.
$H^{(l)}$ $=F^{(l)}(H^{(l-1)}),1\leq l\leq L,$ (1)
where $L$ stands for the number of layers, $H^{(l)}$ is the output of the $l$-th hidden layer $F^{(l)}$ while $H^{(0)}=E$. Next, a prediction layer is adopted to generate the distribution $P$ of the user’s next preferences for each time step $t$:
$P_{t}=Softmax(H^{(L)}_{t}W+b),$ (2)
in which $W\in\mathbb{R}^{d\times|\mathcal{I}|}$, $b\in\mathbb{R}^{|\mathcal{I}|}$ is the learnable parameters. Finally, to train the model, the auto-regressive generative loss are employed as optimization objectives, i.e., left-to-right supervision signals. Combining the cross-entropy loss with such an objective, the loss function can be written as:
$\mathcal{L}(X;\theta)=-\sum_{t=1}^{n-1}logP(y_{t}|X_{\leq t};\theta)=-\sum_{t=1}^{n-1}logP_{t,y_{t}},$ (3)
where $\theta$ is the parameters of the model, $y_{t}=x_{t+1}$ is the item expect to be predicted at step $t$, and $n$ is the input length.
4 AutoSAM: the Proposed Method
4.1 Overview of the Framework
@@
4.2 User Behavior Sampler
@@
$\pi=Softmax(S/\tau)\in\mathbb{R}^{n\times 2},\;A\sim\pi\in\{0,1\}^{n},$ (5)
@@
4.3 Multi-Objective Rewards
Now we discuss how to design our reward, which plays an important role in learning the optimal $\pi$. As mentioned before, traditional SRS often follow the next item prediction task, which can neither update sampler network directly due to the non-differentiable challenge of discrete actions nor control the sample rate. Besides, such single objective may be too strict and one-sided. We argue that the proper decisions should not only focus on the target item but also consider the coherence with preceding context. Along this line, we incorporate the major factors into the reward estimation, including Future Prediction and Sequence Perplexity.
4.3.1 Future Prediction
@@
$min\;J(\theta,\phi)=min\;\mathbb{E}_{X}\mathbb{E}_{A\sim\pi_{\phi}(A|X)}\left[\mathcal{L}(\hat{X};\theta)\right],$ (6)
where $\theta$ and $\phi$ are the parameters of SRS and sampler, respectively. To optimize this objective, we first transform the prediction loss to an equivalent form:
$\mathcal{L}(\hat{X};\theta)$ $=\sum_{t=1}^{m-1}-logP(\hat{y}_{t}|\hat{X}_{\leq t};\theta)$ (7)
@@
$\nabla_{\phi}J(\theta,\phi)=\mathbb{E}_{X}\nabla_{\phi}\sum_{A}\pi_{\phi}(A|X)\mathcal{L}(\hat{X};\theta)$ (8)
@@
where $r^{pre}=\sum_{t=1}^{n-1}-\mathcal{L}_{t}(X,A;\theta)$ means the rewards associated to the sampler. In practice, we may make a few changes to it. First, we add a baseline $\mathcal{L}^{b}$ computed on raw input to $\mathcal{L}$ as it is always positive. Second, we discount the future rewards by a factor $\gamma\in[0,1]$ and ignore the targets whose sample probabilities are lower than $\psi$ to increase stability. Then, the future prediction reward at each time step $t$ can be defined as follows:
$r^{pre}_{t}=\sum_{t^{\prime}=t}^{n-1}-\gamma^{t^{\prime}-t}I[\pi_{\phi}(a_{t^{\prime}+1}|X)>\psi](\mathcal{L}_{t^{\prime}}(X,A;\theta)-\mathcal{L}_{t^{\prime}}^{b}),$ (9)
@@
4.3.2 Sequence Perplexity
In addition to sampling based on future predictions, it should also be attached great importance to considering the coherence with previous context. Inspired by perplexity (PP), which is widely used to evaluate the quality of sentences in Natural Language Processing (NLP) (Liu et al., 2016; Dathathri et al., 2019) with the definition as:
$PP(X)=P(x_{1},x_{2},...x_{n})^{-\frac{1}{n}}=\prod_{t=1}^{n}P(x_{t}|X_{<t})^{-\frac{1}{n}}.$ (10)
As words causing high perplexity tend to stray from the context Lin et al. (2024), similarly, we insist that the items causing high perplexity may lack representativeness. A natural idea is to assign lower sampling probabilities to the items whose predicted probabilities according to historical behaviors are lower than $1/PP(X)$, so as to reduce the perplexity and obtain informative sub-sets in some way. To this end, the sampler should be encouraged to drop $x_{t}$ ($a_{t}=0$) while $P(x_{t}|X_{<t})<1/PP(X)$, and select $x_{t}$ ($a_{t}=1$) while $P(x_{t}|X_{<t})>1/PP(X)$.
In practice, we employ the SRS to compute the approximate objective at each time step $t$. Besides, a relax factor is further conducted to control the strictness of the sampler. Above all, the reward can be defined as follow:
$r^{pp}_{t}=\left(logP(x_{t}|X_{<t})-log\frac{1}{PP(X)}+b\right)(2a_{t}-1),$ (11)